OpenStudy (anonymous):

taylor expansion ln(x+1)

OpenStudy (turingtest):

about x=0 I assume?

OpenStudy (turingtest):

the Taylor series of a function about a point x=a is given by\[f(x)=\sum_{n=0}^\infty{f^{(n)}(a)\over n!}(x-a)^n\]

OpenStudy (turingtest):

if as I assume, you just want it about x=0 you get\[f(x)=\sum_{n=0}^\infty{f^{(n)}(0)\over n!}x^n\]

OpenStudy (anonymous):

true this is a well known series, and you will get it using the method above, but it is usually computed by considering the series \[\frac{1}{1+x}=\sum_{k=0}^{\infty}(-1)^{k+1}x^k=1-x+x^2-x^3+...\]

OpenStudy (anonymous):

since \(\frac{d}{dx}\ln(1+x)=\frac{1}{1+x}\) a simple method is to take the taylor series for \(\frac{1}{1+x}\) and integrate term by term

OpenStudy (turingtest):

so where exactly are you having trouble with this?

OpenStudy (turingtest):

this is the expansion around x=0 right?

OpenStudy (anonymous):

yes, it is!

OpenStudy (turingtest):

\[f(x)=\sum_{n=0}^\infty{f^{(n)}(a)\over n!}(x-a)^n\]you want the expansion about x=0 so this becomes\[f(x)=\sum_{n=0}^\infty{f^{(n)}(0)\over n!}x^n\] so the first term is\[{f(0)\over 0!}x^0\]and what is that equal to?